OpenStudy (anonymous):

Sam has a paper route. Each week she earns $10.00 plus $0.10 for each paper she delivers. Last week she earned $20.00. How many papers did she deliver? Explain your problem solving strategy.

OpenStudy (ghazi):

for each paper she makes 10.10 $ therefore for each dollar she delivers 1/10.10 papers therefore for 20$ she sold 20/10.10 papers

OpenStudy (ghazi):

2 papers

OpenStudy (anonymous):

So she sold 1 paper

OpenStudy (anonymous):

@ghazi incorrect

OpenStudy (ghazi):

see it's 1.98 papers so approximately i took it 2

OpenStudy (anonymous):

she earns $10, regardless of how many papers she sells. If she sells any, she makes $0.10 for each one she sells. So she delivered:\[\frac{20-10}{0.1}=\frac{10}{0.1}=100\]she delivered 100 papers.

OpenStudy (ghazi):

still perplexed

OpenStudy (anonymous):

esedli you are wrong because she only made 20 dollars

OpenStudy (anonymous):

@de4354 So, she delivered 100 papers according to me. She make $0.10 per paper, PLUS $10.00 regardless of her sales. The equates to:\[100($0.10)+$10.00=$20.00\]Where's the problem?

OpenStudy (ghazi):

@eseidl sorry i agree!!

OpenStudy (anonymous):

nevermind i get it now i had to lok at that

OpenStudy (anonymous):

@de4354 You're welcome :)